OpenStudy (anonymous):

Please Help! Solve (sin2A)(tanA)=1 for 0º≤A≤180º. Possible answers: a. 45;135 b. 60;120 c. 30;150 d. 0;180

OpenStudy (aaronq):

use the identity: \(\sf sin(2x) = 2 sin x cos x\) and \(\sf tanx=\dfrac{sinx}{cosx}\) \(\sf (2 sin x cos x)(\dfrac{sinx}{cosx})=1\) \(\sf (2 sin x)(sinx)=1\) \(\sf 2 sin^2x=1\) \(\sf sin^2x=\dfrac{1}{2}\) \(\sf sinx=\sqrt{\dfrac{1}{2}}\) \(\sf x=sin^{-1} (\dfrac{1}{\sqrt2})\) use the unit circle now